# Knolling Bot: Learning Robotic Object Arrangement from Tidy Demonstrations

### v2

@@ Abstract
Addressing the challenge of organizing scattered items in domestic spaces is complicated by the diversity and subjective nature of tidiness. Just as the complexity of human language allows for multiple expressions of the same idea, household tidiness preferences and organizational patterns vary widely, so presetting object locations would limit the adaptability to new objects and environments. Inspired by advancements in natural language processing (NLP), this paper introduces a self-supervised learning framework that allows robots to understand and replicate the concept of tidiness from demonstrations of well-organized layouts, akin to using conversational datasets to train Large Language Models(LLM). We leverage a transformer neural network to predict the placement of subsequent objects. We demonstrate a “knolling” system with a robotic arm and an RGB camera to organize items of varying sizes and quantities on a table. Our method not only trains a generalizable concept of tidiness, enabling the model to provide diverse solutions and adapt to different numbers of objects, but it can also incorporate human preferences to generate customized tidy tables without explicit target positions for each object.
I INTRODUCTION
@@
Designing a robot for household tasks has always presented unique challenges [1, 2, 3, 4]. Unlike industrial settings characterized by uniformity and limited object variety, household environments are filled with distinct objects. Recent progress in object rearrangement has shown robots can organize objects in partially arranged scenes, cluttered tabletops, and constrained environments[5, 6, 7]. While they provided effective approaches to solving task-specific problems, it is crucial to equip the robot with a task-agnostic planner to organize objects based on a generalized sense of tidiness.
@@
The objects in our daily environments, such as desks and houses, vary constantly. In such dynamic settings, providing specific goals for each object is impractical and limits the robot’s generalizability. Instead, robots should be capable of organizing the environment without relying on specific target positions or constant human supervision. This requires the robot to understand and apply a broad concept of organization that extends beyond the specifics of any single task. The necessary adaptability is similar to human-like cognition, where decisions are made not only based on spatial arrangement but also considering the aesthetic and functional aspects of tidiness.
@@
“Knolling” is one such concept intrinsic to humans, referring to the intuitive ability to organize items in a manner that is both aesthetically pleasing and space-efficient (Fig. 1A). Our solution lies in decoupling the cognitive model – which encapsulates the essence of knolling – from the other modules, such as the visual perception system and motor controllers[8]. This division enhances the modularity and interpretability of the entire system compared to only training a single policy mapping from observation to action.
Figure 1: Objects for real-world experiments in a knolling layout. A) A batch of small items, including daily necessities. B) A batch of big items fabricated by a 3d printer.
@@
Given the subjectivity of tidiness and the multiple optimal solutions that could satisfy different individuals’ preferences, knolling is an abstract problem without any definitive standards. Consider the scenario depicted in Fig.2A, where a robotic arm performs a task to integrate a yellow motor into a partially organized environment. The decision on where to place the motor can be depended on various factors, including object category, color, and the need for space efficiency. Training a model solely on regression to predict target positions based on object states might lead to local optimal outcomes. For instance, a model optimized to minimize the loss across different placement preferences could converge to an average of the three potential target positions, resulting in the motor being placed overlapped with a utility knife. This is just an example of a one-step rearrangement of a partially organized scenario. A wrong-placed object may cause the next step of rearrangement to become worse because the errors have accumulated. Moreover, quantifying tidiness with a singular metric is challenging, as no simple equation can encapsulate the diverse aspects of what constitutes a “tidy” space.
In this work, we propose a self-supervised learning framework for modeling tidiness through demonstrations, akin to training a Large Language Model (LLM) with human conversational datasets[9, 10, 11, 12]. We draw parallels between knolling objects and language, where objects act as individual “words” that can be combined in various ways to form “sentences” with identical meanings. Besides, transformer architecture can handle varying input and output sizes with autoregression, which fits the knolling task as the number of objects in the environment is arbitrary[13, 14]. Leveraging the Gaussian Mixture Model (GMM), we address the challenges of multi-label prediction to perform the knolling task[15].
This research aims to contribute to bridging this gap by imbuing a robot with an understanding of the human concept of neatness and developing an integrated pipeline to demonstrate how to deploy it to perform real-world knolling tasks. The pipeline shown in Fig.2B comprises three stages: a knolling model, a visual perception model, and a robot arm controller. The knolling model, based on a transformer architecture, predicts the target positions of objects. The visual perception model, based on our customized YOLO v8, detects objects from an RGB image [16]. Finally, the robot arm controller, leveraging the outputs from the previous stages, guides the robot arm to execute the knolling tasks.
@@
1. We propose a novel self-supervised learning framework for modeling the concept of tidiness. Inspired by advancements in natural language processing, our approach treats objects as “words” that can be organized into aesthetically pleasing and space-efficient “sentences”, allowing the model to learn generalizable tidiness patterns from demonstrations.
2. We leverage transformer architectures and Gaussian Mixture Models to address the multi-label prediction challenges inherent in knolling tasks, where multiple optimal solutions may exist. This enables our model to generate diverse, tidy arrangements tailored to different preferences.
@@
3. We present an integrated end-to-end pipeline that combines the learned knolling model with visual perception and robotic control modules. This allows for deploying knolling capabilities on a real robotic system to organize cluttered environments with varying object quantities and types.
4. Through our experiments, we demonstrate the effectiveness of our approach in generating tidy object layouts, its generalization to unseen objects, and its ability to adapt to user preferences, highlighting the potential for deploying such systems in dynamic household environments.
II RELATED WORK
@@
The domain of robot learning and manipulation tasks has seen significant advancements in recent years. Various manipulation tasks, such as wiping and polishing, stacking, peg-in-hole, and pick-and-place tasks, have been explored[17, 18, 19, 20, 21]. In the realm of stacking tasks, Lee et al. used offline reinforcement learning (RL) to improve upon existing policies for robotic stacking of objects with complex geometry, while Furrer et al. proposed an algorithm for suggesting stable poses for stacking, validated through a real-world autonomous stacking workflow[22, 23]. For pick-and-place tasks, Gualtieri et al. proposed a deep reinforcement learning approach, and Zeng et al. introduced the Transporter Network for vision-based manipulation tasks[24, 25].
@@
In the domain of machine learning models, transformers have revolutionized many fields with their attention-based architecture[26, 27, 28, 29, 30, 31]. The success of transformer models in NLP has prompted exploration in their application to robotic tasks[32, 33, 34]. Jangir et al. proposed the use of transformers with a cross-view attention mechanism for effective fusion of visual information from both cameras for RL policies[35]. Zhu et al. introduced VIOLA, which used a transformer-based policy to improve the robustness of deep imitation learning algorithms[36]. Shridhar et al. proposed PerAct, which uses the preceiver transformer to encode language goals and RGB-D voxel observations[37]. Jain et al. proposed the Transformer Task Planner, which could be pre-trained on multiple preferences, and Dasari et al. explored one-shot visual imitation learning using the Transformer architecture[38].
@@
In the object rearrangement field, most recent works focused on rearranging objects based on explicit instructions[39, 40]. Some other works use a more general method to train a representation model or reward function to supervise the object rearrangement policy without explicit human supervision[41, 42]. LLM progresses, many works leverage the LLM to process natural language instructions from humans and provide robot goal locations for object rearrangement[43, 44, 45]. Instead of giving natural language instructions to the robot, some other works train a robot a command sense of how to place the objects through demonstrations[38, 46, 47]. Combining both approaches, Housekeep[48] provides a method that leverages LLM to train a common sense of tidiness for robots. While these works have contributed significantly to robot manipulation, we aim to provide a common sense of tidiness for robots through demonstrations. We are focusing on knolling tasks that rearrange messy objects on a desktop into a neat layout in the real world. Our approach allows for the handling of varying object types, sizes, and quantities, thus enabling the robot to generate an aesthetically pleasing and space-efficient arrangement of items, similar to human performance.
@@
Figure 3: Knolling model learning framework: The first step in our pipeline involves recognizing and identifying the objects present on a table using the visual perception model. This model takes the input image and outputs n objects’ size (width and length) and position presented as a list. The Knolling Model utilizes a transformer-based model that takes the high-dimensional object state data produced by Positional Encoding as input. The model predicts N target positions after N iterative process.
III METHOD
Our approach of leveraging self-supervised learning on demonstrations of tidy arrangements to train a generalized knolling model is not only effective for desktop object organization but also scalable to broader tidiness tasks. Once a robot can learn the knolling task from tidy desktop demonstrations, the same approach could potentially be extended to larger environments. For example, well-organized room demonstrations could be used to train robots to perform general housekeeping and cleaning tasks. This is analogous to how large language models learn general language understanding from vast datasets of natural conversations and texts.
III-A Data Representation and Generation
We represent objects solely based on their dimensional parameters (width and length), as these are the fundamental properties governing spatial organization. Semantic attributes like color and category are non-differentiable and may introduce subjective biases, so we exclude them from the input data.
To generate a diverse dataset of tidy arrangements, we employ an optimization strategy that iteratively adjusts object placements to minimize the occupied area on the table, subject to different preferences (e.g., color, dimension, category). This stochastic process yields 2.4 million demonstrations spanning a wide variety of tidiness concepts and aesthetic preferences.
For objects with irregular shapes (Fig. 1B), we train a visual perception system to perform segmentation and compute the minimum bounding box, enabling our framework to accommodate unconventional geometries.
III-B Gaussian Mixture Model for Multi-target Learning
$p(S_{i})=\sum_{j=1}^{J}\pi_{ij}\mathcal{N}(S_{i}|\mu_{ij},\sigma_{ij}^{2})$ (1)
III-C Training Knolling Model
We used transformer architecture with encoder-decoder and autoregressive mechanism to iteratively predict object positions based on the object shape. The encoder learns all the shapes of the objects on the table, and the decoder is used to predict the target distributions. The encoder takes a list of the object’s width and length as input $O=\{o_{n}|o_{n}=(w_{n},l_{n}),n\in\mathbb{N}^{+}\}$. The knolling pipeline we presented is shown in Fig. 3.
Leveraging the autoregression of the transformer model, we initialize the model with fully-masked data ($M_{1},...,M_{n}$ in the Fig.3), which is then combined with the encoder’s output to predict the target position of the first object. Once the first object’s position is predicted, it replaces the first element of the masked data, and the model proceeds to predict the target position for the second object. This iterative process continues until all object positions are predicted. The knolling model can be expressed as:
$P_{n}=Dnc\left(Enc(O),\{P_{1},P_{2},\ldots,P_{n-1}\}\right),n\in\mathbb{N}^{+}$
We utilize 2 million data for training and 400k data for testing. The training process leverages a combination of five loss functions to train a model to predict object positions without overlaps while also promoting diversity in predictions:
@@ 2. MSE Loss ($L_{\text{MSE}}$)
The Mean Squared Error (MSE) Loss is an intuitive choice. It computes the squared difference between the sampled predicted position ($\hat{S}$) and the target position ($S$):
@@
where $\hat{S}_{i}$ is the sampled position from the predicted distribution for the $i$-th object, and $S_{i}$ is the actual target position. In a multi-target task, relying solely on MSE could lead the model to predict average positions in the presence of ambiguity, which is not desirable. The MSE Min Loss addresses this by allowing the model to choose the prediction (from among multiple Gaussian components) that is closest to the actual target position.

### v3

# Knolling Bot: Teaching Robots the Human Notion of Tidiness

@@ Abstract
For robots to truly collaborate and assist humans, they must understand not only logic and instructions, but also the subtle emotions, aesthetics, and feelings that define our humanity. Human art and aesthetics are among the most elusive concepts—often difficult even for people to articulate—and without grasping these fundamentals, robots will be unable to help in many spheres of daily life. Consider the long-promised robotic butler: automating domestic chores demands more than motion planning; it requires an internal model of cleanliness and tidiness—a challenge largely unexplored by AI. To bridge this gap, we propose an approach that equips domestic robots to perform simple tidying tasks via knolling, the practice of arranging scattered items into neat, space-efficient layouts. Unlike the uniformity of industrial settings, household environments feature diverse objects and highly subjective notions of tidiness. Drawing inspiration from NLP, we treat knolling as a sequential prediction problem and employ a transformer-based model to forecast each object’s placement. Our method learns a generalizable concept of tidiness, generates diverse solutions adaptable to varying object sets, and incorporates human preferences for personalized arrangements. This work represents a step forward in building robots that internalize human aesthetic sense and can genuinely co-create in our living spaces.
1 Introduction
@@
Designing a robot for household tasks has always presented unique challenges [46, 23, 49, 3]. Unlike industrial settings characterized by uniformity and limited object variety, household environments are filled with diverse objects. Recent progress in object rearrangement has shown that robots can organize objects in partially arranged scenes, cluttered tabletops, and constrained environments[34, 11, 30]. While their work has provided effective approaches to solving task-specific problems, it is crucial to develop methods that can learn a generalized representation of tidiness. Our motivation is to enable robots to understand and apply tidiness concepts that improve with more data, unlike traditional rule-based methods that become increasingly complex with more scenarios.
@@
In common daily environments, including both local environments like a desk, and larger environments like a house, the objects encountered vary constantly. In such dynamic settings, providing specific goals for each object is impractical and limits the robot’s generalizability. Instead, robots should be capable of organizing the environment without relying on specific target positions or constant human supervision. This requires the robot to understand and apply a broad concept of organization that extends beyond the specifics of any single task. The necessary adaptability is similar to human-like cognition, where decisions are made not only based on spatial arrangement but also considering the aesthetic and functional aspects of tidiness.
@@
“Knolling” is one such concept intrinsic to humans, referring to the intuitive ability to organize items in a manner that is both aesthetically pleasing and space-efficient (Fig. 1). Our solution lies in decoupling the cognitive model, which encapsulates the representation of knolling, from the other modules, such as the visual perception system and motor controllers[15, 31]. This division enhances the modularity and interpretability of the entire system compared to only training a single policy mapping from observation to action.
Figure 1: Examples from the knolling task. A) A batch of small items, including daily necessities. B) A batch of big items fabricated by a 3D printer. C) Diverse knolling preferences demonstrated in experiments. The model adapts its tidying behavior based on different preferences: object category (group by function), size (large objects first), and color (group by hue).
@@
Given the subjectivity of tidiness and the multiple optimal solutions that could satisfy different individuals’ preferences, knolling is an abstract problem without any definitive standards. Consider the scenario depicted in Fig.2A, where a robotic arm performs a task to integrate a yellow motor into a partially organized environment. The decision on where to place the motor can be depended on various factors, including object category, color, and the need for space efficiency. Training a model solely on regression to predict target positions based on object states might lead to local optimal outcomes. For instance, in Fig.2A a model optimized to minimize the loss across different placement preferences could converge to an average of the three potential target positions, resulting in the motor being placed in a position that is overlapped with the utility knife. This is just an example of a one-step rearrangement of a partially organized scenario. A wrong-placed object may cause the next step of rearrangement to become worse because the errors have accumulated. Moreover, quantifying tidiness with a singular metric is challenging, as no simple equation can encapsulate the diverse aspects of what constitutes a “tidy" space.
In this work, we propose a self-supervised learning framework for modeling tidiness through demonstrations, akin to training a Large Language Model (LLM) with human conversational datasets[6, 47, 1, 33]. We draw parallels between knolling objects and language, where objects act as individual “words" that can be combined in various ways to form “sentences” with identical meanings. Besides, the transformer architecture can handle varying input and output sizes with autoregression, which fits the knolling task as the number of objects in the environment is arbitrary[42, 12]. Leveraging the Gaussian Mixture Model (GMM), we address the multi-label prediction challenges inherent in the knolling task, where multiple valid placements may exist for the same object depending on contextual preferences (e.g., spatial efficiency or category)[37].
Our research contributes by equipping robots with a conceptual understanding of tidiness, demonstrating an integrated pipeline for knolling in real-world tasks. As shown in Fig.2B , itcomprises three stages: a knolling model, a visual perception model, and a robot arm controller. The knolling model, based on a transformer architecture, predicts the target positions of objects. The visual perception model, based on our customized YOLO v8, detects objects from an RGB image [35]. Finally, the robot arm controller, leveraging the outputs from the previous stages, guides the robot arm to execute the knolling tasks.
@@
  1. 1.
    We propose a novel self-supervised learning framework for modeling the representation of tidiness. This allows the model to learn generalizable tidiness patterns directly from demonstrations, improving its performance as more data becomes available.
  2. 2.
    We leverage transformer architectures and Gaussian Mixture Models to address the multi-label prediction challenges inherent in knolling tasks, and can process unseen objects and preferences, highlighting the potential for deploying such systems in dynamic household environments.
  3. 3.
@@
    We demonstrate an integrated pipeline combining the learned knolling model with visual perception and robotic control modules. This allows for deploying knolling capabilities on a real robotic system to organize cluttered environments with varying object quantities and types.
  4. 4.
    We contribute to the research community by not only open-sourcing our dataset of tidy object arrangements but also establishing a benchmark with comprehensive evaluation metrics for the knolling task. This benchmark will enable comparative studies and further exploration of arrangement tasks with arbitrary object numbers and shapes.
2 RELATED WORK
@@
The domain of robot learning and manipulation tasks has seen significant advancements in recent years. Various manipulation tasks, such as wiping and polishing, stacking, peg-in-hole, and pick-and-place tasks, have been explored[4, 51, 26, 24]. In the realm of stacking tasks, Lee et al. used offline reinforcement learning (RL) to improve upon existing policies for robotic stacking of objects with complex geometry, while Furrer et al. proposed an algorithm for suggesting stable poses for stacking, validated through a real-world autonomous stacking workflow[25, 10]. For pick-and-place tasks, Gualtieri et al. proposed a deep reinforcement learning approach, and Zeng et al. introduced the Transporter Network for vision-based manipulation tasks[39, 48].
@@
In the domain of machine learning models, transformers have revolutionized many fields with their attention-based architecture[32, 21, 29, 19, 44, 28]. The success of transformer models in NLP has prompted exploration in their application to robotic tasks[22, 8, 36]. Jangir et al. proposed the use of transformers with a cross-view attention mechanism for effective fusion of visual information from two cameras for RL policies[16]. Zhu et al. introduced VIOLA, which used a transformer-based policy to improve the robustness of deep imitation learning algorithms[50]. Shridhar et al. proposed PerAct, which used the preceiver transformer to encode language goals and RGB-D voxel observations[40]. Jain et al. proposed the Transformer Task Planner, which could be pre-trained on multiple preferences, and Dasari et al. explored one-shot visual imitation learning using the transformer architecture[14].
@@
In the object rearrangement field, most recent work has focused on rearranging objects based on explicit instructions[13, 7]. Some other works use a more general method to train a representation model or reward function to supervise the object rearrangement policy without explicit human supervision[45, 43]. With the rapid rise of LLM as powerful all-purpose models, many works have leveraged them to processs[27, 5, 2]. Instead of giving natural language instructions to the robot, some other works train a robot a common sense of how to place the objects through demonstrations[14, 18, 38]. Combining both approaches, Housekeep[17] provides a method that leverages LLM to train a common sense of tidiness for robots. While these works have contributed significantly to robot manipulation, we aim to provide a common sense of tidiness for robots through demonstrations. We focus on knolling tasks that rearrange messy objects on a desktop into a neat layout in the real world. Our approach allows for the handling of varying object types, sizes, and quantities, thus enabling the robot to generate an aesthetically pleasing and space-efficient arrangement of items, similar to human performance.
@@
Figure 3: Knolling Model Learning Framework: The pipeline begins with the visual perception model, which processes an input image to identify objects and extract their state representations—including width (w), length (l), position and orientation, presented as a list. However, only w and l are used as input for the knolling. The knolling model takes high-dimensional object states (h) derived from positional encoding as input. During training, a masked learning approach is employed, where part of the object data (M) is masked, and the model learns to predict the next object’s position (P). The model predicts N target positions after N iterative processes.
3 METHOD
3.1 Data Representation and Generation
We represent objects solely based on their dimensional parameters (width and length), as these are the fundamental properties governing spatial organization. semantic attributes like color and category are excluded from the model input due to the subjective variability and potential bias they introduce in the context of tidiness. Unlike quantifiable features (e.g., position, size), attributes like color and category are often interpreted differently based on context and individual perception. Focusing on spatial and geometric attributes is more objectively measurable. This approach helps the model generalize tidiness based on universally applicable metrics rather than subjective semantic interpretations that might vary across different settings. To generate a diverse dataset of tidy arrangements, we design an optimization strategy that iteratively adjusts object placements to minimize the occupied area on the table. This strategy can control the results by adjusting placements based on object attributes such as color, dimension, or category, yielding distinct organizational patterns and aesthetic preferences. By iterating through multiple configurations, this stochastic process produces 2.4 million demonstrations that span a broad spectrum of tidiness concepts and visual styles, providing a rich dataset for training. For objects with irregular shapes (Fig. 1B), we train a visual perception system to perform segmentation and compute the minimum bounding box, enabling our framework to accommodate unconventional geometries.
3.2 Training The Knolling Model
Our self-supervised learning framework for training knolling models follows a curriculum-based approach, progressing from simple to complex tasks as shown in Fig. 3. We designed a two-stage training process to first acquaint the model with the basics of object arrangement and then refine its proficiency in executing complete knolling tasks from scratch. By training a representation of tidiness through self-supervised learning, our model’s performance improves as more data becomes available, unlike traditional rule-based methods that become increasingly complex with more scenarios. For further details on the two training phases, pre-training with self-supervised Learning, amd finetuning, please see Appendix L. Our knolling pipeline comprises three modules: a knolling model, a visual perception model, and the robot arm controller for executing the pick-and-place task. For further details on the pipeline and training objectives, see Appendix K.
4 Experiments
4.1 Qualitative Evaluations in the Simulation
Handling Varying Object Quantities. In real-world scenarios, the number of objects on a surface can vary significantly. To assess our model’s ability to handle this variability, we conducted experiments with diverse object counts ranging from 2 to 10 items. Our model, leveraging its autoregressive nature, successfully generated tidy arrangements across all object quantities, as shown in Fig. 4. This adaptability to varying input sizes is a crucial aspect of our approach, enabling it to generalize to different scenarios without explicit retraining.
Generating Diverse Solutions Based on Preferences. Tidiness preferences can differ significantly among individuals, leading to multiple valid solutions for the same set of objects. To capture this diversity, our model decouples preferences from the input data (which consists solely of object dimensions) and instead encodes preferences through the order of the input sequence. By altering the input order based on criteria such as object category, color, or size, our model generates distinct tidy arrangements tailored to specific preferences, as illustrated in Fig. 1 C. This flexibility to incorporate user preferences without modifying the model architecture is a key strength of our approach.
4.2 Real-World Knolling Experiments
To demonstrate the practical applicability of our framework, we deployed our trained knolling model on a 5-DoF robotic arm (WidowX 200) equipped with an Intel Realsense D435 camera for visual perception. We randomly placed 6-10 boxes of varying sizes and colors within the robot’s workspace, simulating cluttered tabletop scenarios. The integrated pipeline (Fig. 2) first captures an overhead image of the cluttered scene, which is processed by our visual perception module to detect objects, perform segmentation, and extract their dimensions and poses. The knolling model then predicts the target positions for each object based on their dimensions. Finally, the robotic arm controller executes smooth pick-and-place operations, guided by the predicted target positions and current object poses, to realize the tidy arrangements. Fig. 5A showcases real-world knolling tasks performed on different box configurations, demonstrating our system’s capability to organize diverse object sets. Fig. 5B further illustrates the system’s adaptability to varying object quantities, successfully knolling scenes with 6, 8, and 10 objects. Additionally, Fig. 5C highlights the generation of distinct tidy arrangements for the same set of objects, reflecting different user preferences based on category, color, or shape. We encourage readers to view our supplementary video for a comprehensive visualization of these real-world knolling experiments. For further quantitative evaluation and an ablation study, please see Appendix I. For an analysis of how performance scales with data, see Appendix J.
5 CONCLUSIONS
In this work, we introduced a self-supervised learning framework for enabling robots to understand and replicate the representation of tidiness, or “knolling," from demonstrations of well-organized object arrangements. Our approach leverages transformer architectures and Mixture Density Networks to model the inherent multi-modality of knolling tasks, where identical object configurations can lead to diverse but tidy arrangements. Our model learns a generalizable representation of tidiness that can adapt to varying object quantities and incorporate user preferences by training on a large dataset of knolling demonstrations without explicit human supervision or target positions. We demonstrated the effectiveness of our framework in generating aesthetically pleasing and space-efficient object arrangements. Our quantitative evaluations showed that our transformer-based model consistently outperformed baseline architectures, highlighting the advantages of self-attention and autoregressive mechanisms for handling variable input sizes and multi-label predictions. Moreover, we presented an integrated pipeline that combines our learned knolling model with visual perception and robotic control modules, enabling the deployment of knolling capabilities on a real robotic system. Our real-world experiments showcased the system’s ability to organize cluttered environments with diverse object types and quantities, adapting to different user preferences based on criteria such as object category, color, or shape. This work represents an important step forward in imbuing robots with human aesthetics, a critical requirement for effective human-robot interaction, and we hope our work inspires further research in this area.
Appendix A Training Objectives
We employ a combination of loss functions during training:
@@ 2. MSE Loss ($L_{\text{MSE}}$)
The Mean Squared Error (MSE) Loss computes the squared difference between the sampled predicted position ($\hat{S}$) and the target position ($S$):
@@
where $\hat{S}_{i}$ is the sampled position from the predicted distribution for the $i$-th object, and $S_{i}$ is the actual target position.